# On the Identification of Exotic Compact Binaries with Gravitational Waves: a Phenomenological approach

### v1

@@ Abstract
Gravitational wave (GW) astronomy has been hailed as a gateway to discovering unexpected phenomena in the universe. Over the last decade there have been close to one hundred GW observations of compact-binary mergers. While these signals are largely consistent with mergers of binary black holes, binary neutron stars, or black hole-neutron star systems, some events suggest the intriguing possibility of binaries involving exotic compact objects (ECOs). Identifying and characterising an ECO merger would require accurate ECO waveform models. Using large numbers of numerical relativity simulations to develop customised models for ECO mergers akin to those used for binary black holes, would be not only computationally expensive but also challenging due to the limited understanding of the underlying physics. Alternatively, key physical imprints of the ECO on the inspiral or merger could in principle be incorporated phenomenologically into waveform models, sufficient to quantify generic properties. In this work we present a first application of this idea to assess the detectability and distinguishability of ECO mergers, and we propose a phenomenological approach that can iteratively incorporate features of ECO mergers, laying the groundwork for an effective exotic compact object identifier in compact binary coalescences. We demonstrate that within this framework the compactness of the objects in GW150914 are consistent with that of black holes. The efficacy of the identifier can be refined by adding information from numerical relativity simulations involving fundamental fields. Conversely, such an identifier framework can help focus future numerical relativity and modeling efforts for exotic objects.
@@ I Introduction
The morphology of the GW signal from an ECO binary is expected to differ from that of a BBH primarily due to differences in compactness of the objects and tidal interactions. This suggests that a careful consideration of the contribution to the phase of the binary due to tidal deformation by a companion, tidal heating or spin-induced quadrupole moments can distinguish ECO binaries from BBHs. In fact the distinguishability of ECO binaries from BBHs has been explored in some detail by perturbative modifications to the inspiral phase [18, 19, 20, 21, 22, 23, 24, 25, 26, 27]. More recently a templated search with an inspiral only model for compact objects was conducted on the publicly available data from the LIGO-Virgo-KAGRA (LVK) collaboration [28] by incorporating effects of large tidal deformabilities into the phase modulation and obtaining model independent constraints on the Love numbers. Additionally, the remnant of a merger of an ECO binary may produce an object of similar origin, or a BH, making ringdown tests a good method for distinguishability have also been considered [29, 30, 31, 32].
@@
To make an approximate phenomenological model based on universal features of the inspiral, we note that compact objects larger than black holes can be expected to merge at a larger separation, and so the inspiral will terminate at a lower frequency. Based on this simple observation we modify the amplitude of a standard BBH waveform model,   PhenomD, to terminate at a lower frequency, parameterised by the object’s compactness. For this first illustrative exercise we do not attempt to model any other physics, e.g., tidal deformations during the inspiral, or the nature of the merged object; we make only one change to the waveform, and assess its impact. Furthermore, a distinguishability criterion for putative BBH detections has not been defined to date. Our aim in this work is to explore the plausible range of physical parameters for ECOs, assess their detectability by current generation GW detectors and propose an “ECO identifier” that can discriminate between a BBH and non-BBH hypothesis, without loss of generality. We note here that a similar early end of inspiral has been noted for neutron star-black hole binaries [42] as well as for exotic binaries [43, 28] based on an earlier peak of the phase derivative due to the inclusion of tidal terms, and have been used in state-of-the-art CBC waveforms that include matter effects [44, 45]. Thus, our amplitude modification does in fact include the most dominant impact of phase modulation due to tidal effects, without making specific assumptions about the underlying physics.
@@ II.2 Tapered model
$150M_{\odot}$ $75M_{\odot}$ $75M_{\odot}$ $58.66$ Hz $29.33$ Hz 125.11 Hz
@@
We use a $tanh$ window to ensure a smooth fall off and the window width, $f_{D}$, gives us some flexibility in adjusting the slope of the fall. We choose a conservative window width of $f_{D}=10$ Hz, as shown in figure. 2 instead of a , for inspecting detectability of ECOs by inspiral-merger-ringdow (IMR) waveforms.
@@ II.3 PhenomDECO
Figure 3: Same as fig. 2 with the amplitude parameterised by compactness. We show the amplitude four different compactness parameter, going from lower to higher values represented by progressively thicker lines. We also show that for C=0.5  PhenomDECO  reproduces the BBH amplitude as shown by the black dashed line.
@@ III Detecting exotic mergers
We explore the detectability of ECO mergers with signals generated using  PhenomDTaper  as well as  PhenomDECO  but show the results only for  PhenomDTaper  because it yields the more conservative estimates between the two. This can be easily understood by the sharp amplitude fall off in PhenomDTaper  that reduces the power in the signal compared to  PhenomDECO . We will focus on objects whose compactness lies in the range $0.50.1\leq C<$ 0.5; $C=0.25$ corresponds to objects twice as large as a Schwarzschild BH of the same mass. We assume the same compactness for both objects and that they are non-spinning. We will consider Bayesian inference to detect ECOs by PhenomDECO  in Sec IV.
@@ III.1 Detecting ECO mergers with a BBH model
The match is computed using the implementation in the pycbc library [55], using the advanced-LIGO detector sensitivity curve [56] with a lower frequency cut-off of 20 Hz, and then we optimize over the intrinsic parameters $(q,M,\chi_{\rm{eff}})$ of the template to obtain the fitting factor,$\mathfrak{F}$ [57].
Figure 5: Fitting factor contours obtained by optimizing the match between PhenomDTaper and  PhenomD over the $q-M-\chi_{\rm{eff}}$ parameter space for different ECO total masses (x-axis) and mass ratio (y-axis); top left: $C=0.1$, top right: $C=0.25$, bottom left: $C=0.33$, bottom right: $C=0.45$ and colour bar on right shows the value of the fitting factors. The black contours on the top left panel enclose the region of $q-M$ space for non-spinning ECOs that would be interesting in residual tests, if such ECOs exist in nature.
@@
For moderately compact objects ($C\gtrsim 0.25$) the $\mathfrak{F}$ never falls below $\sim 0.9$ as long as the binary is not very highly asymmetric in mass. However, the actual ability to detect in a search is a more nuanced discussion. For instance, a $\mathfrak{F}<0.5$ as can be seen for an ECO binary with $C=0.1$ and $M_{\rm{ECO}}=50M_{\odot}$ would not be detectable if the true SNR of the signal is 15, assuming a $\rho_{\rm{threshold}}=8$. However, if the signal has an SNR of 20 or more it would be detected, provided it passes the usual consistency checks done in a BBH search. Nevertheless, fig. 5 gives an indication of the fraction of ECO binaries that would be missed in a population using a BBH search. Noting that the SNR roughly scales as $\rho\propto\frac{1}{D_{L}}$, $D_{L}$ being luminosity distance to the source, a waveform that recovers $0.9\rho$ is sensitive to only a distance of $0.9D_{L}$ (and a volume of $0.730.9^{3}=$ 0.73).
@@
[…]
@@ III.2 Distinguishability
Inverting the above equation, we can compute the SNR at which an ECO merger would be distinguishable from a BBH merger. Figure 6 shows the distinguishability SNR for different compactness as function of ECO total mass. It is easy to note from this plot that for low mass systems ($<10M_{\odot}$) and moderate to large compactnesses, it would be almost impossible to distinguish between an ECO merger and a BBH merger. On the other hand for masses $>10M_{\odot}$ distinguishability requires only moderately high SNR for $C\lesssim 0.33$. In fact at masses $\gtrsim 30M_{\odot}$ the distinguishability SNR is always lower than the detectability threshold of 8, suggesting that detected events will always be distinct. It is worthwhile to draw attention to an important caveat of our analysis at this point again; we ignore phase modulations due to finite-sized effects as well as tidal disruptions, and using a model specifically calibrated to such dephasing may significantly change distinguishability. Unfortunately this also means that using a model finely tuned to a particular model of ECO binary is as precarious as using a BBH model, in terms of distinguishability. In other words, unless a generalized framework is used to look for exotic objects, we are more likely to find what we are looking for! We discuss our first attempt in this direction in section IV.
@@ III.3 Source parameters
Figure 7 gives an insight into the extent of biases in the mass parameters. For equal mass ECO signals of total mass $\gtrsim 30$, the template prefers a higher total mass and mass ratio $q_{\rm{BBH}}>1$. To understand the interplay between mass ratio and total mass we again refer to the top panels of Fig. 4. While a higher total mass of the template captures the early truncation of the amplitude, it also increases the overall scaling of the amplitude envelope. The template tries to compensate for this by preferring a higher mass ratio, hence the $M-q$ biases are correlated.
@@ IV.1 Compactness
As shown in Fig. 3, the frequency domain amplitude of PhenomDECO  approaches that of a BBH merger for $C=0.5$. Therefore, we first want to verify that compactness estimates obtained using  PhenomDECO  are reliable. We generate a BBH signal with a total mass of $40M_{\odot}$ at $q=1$ and $\chi_{\rm{eff}}=0.1$ using PhenomDECO  and inject it at an SNR of $\sim 25$ into stationary Gaussian noise at design sensitivity of the advanced LIGO detectors [56] with zero detector noise. We attempt to recover it with  PhenomDECO  by sampling over compactness in addition to the intrinsic parameters of chirp mass ($\mathcal{M}_{c}$), mass ratio ($q$), aligned spin components ($\chi_{1,2}$), coalescence phase ($\phi_{c}$), and the extrinsic parameters of distance ($d_{L}$), sky location ($\alpha,\delta$), polarization ($\psi$) & inclination ($\iota$). To ensure robustness of the compactness estimate we assigned a uniform distribution of prior probability in the range $\in[0.1,0.99]$. Note that, a compactness $>0.5$ is indicative of a Kerr BH.
@@ IV.2 Compactness estimates of real GW events
While the injections above demonstrate the significance compactness has on the signal, the robustness of the compactness measure can only be verified by estimates from real GW events. We performed Bayesian parameter estimation on GW150914 [62] to infer compactness. The center panel on fig. 10 shows that PhenomDECO is able to 1) find a clear peak for the posterior probability distribution of compactness and 2) constrain $95\%$ of posterior probability density for compactness (i.e., $95\%$ credible interval) within $[0.505,0.616]$. Nominally, a compactness of $C=0.5$ is excluded at $95\%CI$, however, GW150914 has some support for aligned spin content as well as mass asymmetry in the binary as shown by the left and right panels of fig. 10 respectively. We know that a Kerr BH has a higher compactness than a Schwarzschild BH, although the two-dimensional posterior in Fig. 11 does not show a strong correlation. This figure also reveals that while the 1D projection of the joint probability distribution excludes 0.5 at 95% CI, it is actually included in the $C-\chi_{\rm{eff}}$ two-dimensional marginalized posterior probability distribution.
@@ V Conclusions
We find that for objects with $C\gtrsim 0.25$, i.e., up to twice as large as black holes, current BBH searches would rarely fail to detect observable signals for masses below 10 $M_{\odot}$. Only for much lower compactness (i.e., objects 5 times larger than black holes) would we fail to detect most binaries with masses above 10 $M_{\odot}$. However, the statement on detectability is incomplete without consideration of SNR distribution of the sources. Notably, the results in fig. 5 can also be interpreted as the fractional loss in SNR in an observation, so that a merger signal of two $C=0.1$ ECOs and total mass of $30M_{\odot}$ with loud enough SNR (say 100) would be easily detectable by a  PhenomD  template, recovering an SNR50similar-toabsent50$\sim$ 50∼ 50.
@@
We also consider the parameter biases for ECOs measured assuming they are BBH. We find that the mass ratio is biassed towards more asymmetric masses. This suggests that, when considering whether any current observations are ECOs, we should focus on those measured to have unequal masses. For instance, an event like GW190814 [63], where we observed a high mass ratio, could be an ECO, but based on fig. 5 must have $C>0.33$ since that the best-fitting BBH template has a fitting factor of $0.97$[59].
@@
These measurement results highlight a number of caveats to this work. We have included only one physical effect in potential ECOs, namely a simple model of the effect of compactness on the contact point that terminates inspiral. Not all ECOs will follow this phenomenology. For example, consider one of the boson star simulations in Ref. [39] where, even with a compactness of $C=0.2$, the signal is almost identical to that from a BBH. Aside from ignoring tidal effects and post-merger dynamics, this is also partly because that work uses a different interpretation for compactness than the usual $M/R$ definition used for BHs. This reminds us that our results apply to ECOs where compactness impacts the signal in one specific way, but does not necessarily apply to all ECOs, or all definitions of compactness.
@@
Similarly, a strength of such an approach also is that it allows us to determine the properties of compact objects that could be measured without making any prior assumptions about the nature of the object. By contrast, if one produces a model of signals from, for example, boson-star binaries, and then observes a signal consistent with that model (and not consistent with BBH), we cannot then conclude that the objects are in fact boson stars: only that they are more like boson stars than black holes. They may in fact be the manifestation of some entirely different physical phenomena that leads to compact objects that, when in binaries, produce GW signals more akin to boson-star binaries than black-hole binaries. A more agnostic approach would allow us to measure general properties of ECOs, and from there determine the range of physical processes that could produce such objects. Once we identify which physical properties have the largest influence on the observed signal morphology, we can prioritize these features in follow-up NR simulations. Therefore, compactness estimates from PhenomDECO can inform NR simulation priorities – and of course, simulation results are necessary to refine PhenomDECO  as an ECO identifier.

### v3

@@ Abstract
Gravitational wave (GW) astronomy has been hailed as a gateway to discovering unexpected phenomena in the universe. Over the last decade there have been close to one hundred GW observations of compact-binary mergers. While these signals are largely consistent with mergers of binary black holes, binary neutron stars, or black hole-neutron star systems, some events suggest the intriguing possibility of binaries involving exotic compact objects (ECOs). Identifying and characterising an ECO merger would require accurate ECO waveform models. Using large numbers of numerical relativity simulations to develop customised models for ECO mergers akin to those used for binary black holes, would be not only computationally expensive but also challenging due to the limited understanding of the underlying physics. Alternatively, key physical imprints of the ECO on the inspiral or merger could in principle be incorporated phenomenologically into waveform models, sufficient to quantify generic properties. In this work we present a first application of this idea to assess the detectability and distinguishability of ECO mergers, and we propose a phenomenological approach that can iteratively incorporate features of ECO mergers, laying the groundwork for an effective exotic compact object identifier in compact binary coalescences. Using Bayesian parameter estimation on the data for the GW event GW150914, we find the inferred compactness to be consistent with that expected for black holes, within this framework. The efficacy of the identifier can be refined by adding information from numerical relativity simulations involving fundamental fields. Conversely, such an identifier framework can help focus future numerical relativity and modeling efforts for exotic objects.
@@ I Introduction
The morphology of the GW signal from an ECO binary is expected to differ from that of a BBH primarily due to differences in compactness of the objects and tidal interactions. This suggests that a careful consideration of the contribution to the phase of the binary due to tidal deformation by a companion, tidal heating, spin-induced quadrupole moments or self-interactions, in case of boson stars, can distinguish ECO binaries from BBHs. In fact the distinguishability of ECO binaries from BBHs has been explored in some detail by perturbative modifications to the inspiral phase [18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30]. More recently a templated search with an inspiral only model for compact objects was conducted on the publicly available data from the LIGO-Virgo-KAGRA (LVK) collaboration [31] by incorporating effects of large tidal deformabilities into the phase modulation and obtaining model independent constraints on the Love numbers. Additionally, the remnant of a merger of an ECO binary may produce an object of similar origin, or a BH, making ringdown tests a good method for distinguishability have also been considered [32, 33, 34, 35].
@@
To make an approximate phenomenological model based on universal features of the inspiral, we note that compact objects larger than black holes can be expected to merge at a larger separation, and so the inspiral will terminate at a lower frequency. Based on this simple observation we modify the amplitude of a standard BBH waveform model,   PhenomD, to terminate at a lower frequency, parameterised by the object’s compactness. For this first illustrative exercise we do not attempt to model any other physics, e.g., tidal deformations during the inspiral, or the nature of the merged object; we make only one change to the waveform, and assess its impact. Furthermore, a distinguishability criterion for putative BBH detections has not been defined to date. Our aim in this work is to explore the plausible range of physical parameters for ECOs, assess their detectability by current generation GW detectors and propose an “ECO identifier” that can discriminate between a BBH and non-BBH hypothesis, without loss of generality. We note here that a similar early end of inspiral has been noted for neutron star-black hole binaries [45] or sub-solar mass binaries [46], as well as for exotic binaries [47, 31], based on an earlier peak of the phase derivative due to the inclusion of tidal terms, and have been used in state-of-the-art CBC waveforms that include matter effects [48, 49]. Thus, our amplitude modification does in fact include the most dominant impact of phase modulation due to tidal effects, without making specific assumptions about the underlying physics.
@@ II.2 Tapered model
$150M_{\odot}$ $_{\odot} 75M75M_{\odot}$ $58.66$ Hz $29.33$ Hz 125.11 Hz
@@
We use a tanhtanh window to ensure a smooth fall off and the window width, $f_{D}$, gives us some flexibility in adjusting the slope of the fall. We choose a conservative window width of $f_{D}=10$ Hz, as shown in figure. 2 instead of a , for inspecting detectability of ECOs by inspiral-merger-ringdow (IMR) waveforms.
@@ II.3 PhenomDECO
Figure 3: Same as fig. 2 with the amplitude parameterised by compactness. We show the amplitude for four different values of the compactness parameter, going from lower to higher values represented by progressively thicker lines. We also show that for C=0.5  PhenomDECO  reproduces the BBH amplitude as shown by the black dashed line.
@@ III Detecting exotic mergers
We explore the detectability of ECO mergers with signals generated using  PhenomDTaper  as well as  PhenomDECO  but show the results only for  PhenomDTaper  because it yields the more conservative estimates between the two. This can be easily understood by the sharp amplitude fall off in PhenomDTaper  that reduces the power in the signal compared to  PhenomDECO . We will focus on objects whose compactness lies in the range $0.1\leq C<0.5$; $C=0.25$ corresponds to objects twice as large as a Schwarzschild BH of the same mass. We assume the same compactness for both objects and that they are non-spinning. We will consider Bayesian inference to detect ECOs by PhenomDECO  in Sec IV.
@@ III.1 Detecting ECO mergers with a BBH model
$\langle h_{\rm{PhenomD}}|h_{\rm{ecomodel}}\rangle=4\mathfrak{R}\int\frac{h_{\rm{PhenomD}}(f)h_{\rm{ecomodel}}^{*}(f)}{S_{n}(f)}df.$
@@
Figure 5: Fitting factor contours obtained by optimizing the match between PhenomDTaper and  PhenomD over the $q-M-\chi_{\rm{eff}}$ parameter space for different ECO total masses (x-axis) and mass ratio (y-axis); top left: $C=0.1$, top right: $C=0.25$, bottom left: $C=0.33$, bottom right: $C=0.45$ and colour bar on right shows the value of the fitting factors. The black contours on the top left panel enclose the region of qMq-M space for non-spinning ECOs that would be interesting in residual tests, if such ECOs exist in nature.
@@
For moderately compact objects ($C\gtrsim 0.25$) the $\mathfrak{F}$ never falls below $\sim 0.9$ as long as the binary is not very highly asymmetric in mass. However, the actual ability to detect in a search is a more nuanced discussion. For instance, a $\mathfrak{F}<0.5$ as can be seen for an ECO binary with $C=0.1$ and $M_{\rm{ECO}}=50M_{\odot}$ would not be detectable if the true SNR of the signal is 15, assuming a $\rho_{\rm{threshold}}=8$. However, if the signal has an SNR of 20 or more it would be detected, provided it passes the usual consistency checks done in a BBH search. Nevertheless, fig. 5 gives an indication of the fraction of ECO binaries that would be missed in a population using a BBH search. Noting that the SNR roughly scales as $\rho\propto\frac{1}{D_{L}}$, $D_{L}$ being luminosity distance to the source, a waveform that recovers $0.9\rho$ is sensitive to only a distance of $0.9D_{L}$ (and a volume of $0.9^{3}=0.73$).
@@
For $C=0.1$ panel in Fig. 5 the black contour line encloses the region that will lead to sufficiently high residual SNR to indicate an exotic origin for the merger signal provided the signal has at least an SNR of $\sqrt{8^{2}+\rho_{\rm{BBH}}^{2}}$. Consistent with expectations, lower values of recovered $\rho_{\rm{BBH}}$ for a similar residual would make the test sensitive to a wider region of ECO mass in the $q_{\rm{ECO}}$$M_{\rm{ECO}}$ parameter space. This can be easily spotted in table  2 where we have shown a selection of events from the O3 catalog [4] for which residual test results were reported [63]. While GW190519_153544 and GW190828_065509 have similar residual SNR, the event with the lower SNR has significantly lower $\mathfrak{F}$. On the other hand GW190706_222641 and GW190708_232457 have similar recovered SNR, but significantly different residual SNR, with the higher one corresponding to a lower $\mathfrak{F}$. The residuals for all of the events reported in [63] were found to be consistent with noise, however, it is worth noting that this test is currently performed with the motivation of finding deviations from GR. Therefore, unless recovered SNRs and residuals show a correlation across all events, residuals in individual events are ruled out as noise artefacts. On the other hand if the catalogs comprise true BBH mergers and a set of ECO mergers mischaracterised as BBH mergers, we would expect only a few to have significant residuals at arbitrary recovered SNRs. Given the significantly high residual left behind by GW190828_065509 despite the low recovered SNR (in comparison to other events of similar residual), it may be interesting to test further with exotic waveforms.
@@ III.2 Distinguishability
Inverting the above equation, we can compute the SNR at which an ECO merger would be distinguishable from a BBH merger. Figure 6 shows the distinguishability SNR for different compactness as function of ECO total mass. It is easy to note from this plot that for low mass systems ($10M_{\odot}$) and moderate to large compactnesses, it would be almost impossible to distinguish between an ECO merger and a BBH merger. On the other hand for masses $10M_{\odot}$ distinguishability requires only moderately high SNR for $C\lesssim 0.33$. In fact at masses $\gtrsim 30M_{\odot}$ the distinguishability SNR is always lower than the detectability threshold of 8, suggesting that detected events will always be distinct. It is worthwhile to draw attention to an important caveat of our analysis at this point again; we ignore phase modulations due to finite-sized effects as well as tidal disruptions, and using a model specifically calibrated to such dephasing may significantly change distinguishability. Unfortunately this also means that using a model finely tuned to a particular model of ECO binary is as precarious as using a BBH model, in terms of distinguishability. In other words, unless a generalized framework is used to look for exotic objects, we are more likely to find what we are looking for! We discuss our first attempt in this direction in section IV.
@@ III.3 Source parameters
Figure 7 gives an insight into the extent of biases in the mass parameters. For equal mass ECO signals of total mass $\gtrsim 30$, the template prefers a higher total mass and mass ratio $q_{\rm{BBH}}>1$. To understand the interplay between mass ratio and total mass we again refer to the top panels of Fig. 4. While a higher total mass of the template captures the early truncation of the amplitude, it also increases the overall scaling of the amplitude envelope. The template tries to compensate for this by preferring a higher mass ratio, hence the MqM-q biases are correlated.
@@ IV An ECO identifier
Being built upon  PhenomD ,  PhenomDECO naturally inherits its regime of validity in generating reliable waveforms generation. Specifically, configurations beyond $q=18$ and $\chi_{\rm{eff}}=0.85$ (0.98 for equal mass binaries) fall outside calibration region and the model should be used with caution in this region. Additionally, we have tested the model in the range $C>0.05$ and $q<20$, and do not see signs of pathological behaviour. The present version of the model is not valid beyond the Kerr limit.
@@ IV.1 Compactness
As shown in Fig. 3, the frequency domain amplitude of PhenomDECO  approaches that of a BBH merger for $C=0.5$. Therefore, we first want to verify that compactness estimates obtained using  PhenomDECO  are reliable. We generate a BBH signal with a total mass of $40M_{\odot}$ at $q=1$ and $\chi_{\rm{eff}}=0.1$ using PhenomDECO  and inject it at an SNR of $\sim 25$ into stationary Gaussian noise at design sensitivity of the advanced LIGO detectors [60] with zero detector noise. We attempt to recover it with  PhenomDECO  by sampling over compactness in addition to the intrinsic parameters of chirp mass ($\mathcal{M}_{c}$), mass ratio ($q$), aligned spin components ($\chi_{1,2}$), coalescence phase ($\phi_{c}$), and the extrinsic parameters of distance ($d_{L}$), sky location ($\alpha,\delta$), polarization ($\psi$) & inclination ($\iota$). To ensure robustness of the compactness estimate we assigned a uniform distribution of prior probability in the range $\in[0.1,0.99]$. Note that, a compactness >0.5>0.5 is indicative of a Kerr BH.
@@ IV.2 Compactness estimates of real GW events
While the injections above demonstrate the significance compactness has on the signal, the robustness of the compactness measure can only be verified by estimates from real GW events. We performed Bayesian parameter estimation on the LIGO Hanford and Livingston data from the GW150914 observation [66] to infer compactness. The center panel on fig. 10 shows that PhenomDECO is able to 1) find a clear peak for the posterior probability distribution of compactness and 2) constrain $95\%$ of posterior probability density for compactness (i.e., $95\%$ credible interval) within $[0.505,0.616]$. Nominally, a compactness of $C=0.5$ is excluded at $95\%CI$, however, GW150914 has some support for aligned spin content as well as mass asymmetry in the binary as shown by the left and right panels of fig. 10 respectively. We know that a Kerr BH has a higher compactness than a Schwarzschild BH, although the two-dimensional posterior in Fig. 11 does not show a strong correlation. This figure also reveals that while the 1D projection of the joint probability distribution excludes 0.5 at 95% CI, it is actually included in the $C-\chi_{\rm{eff}}$ two-dimensional marginalized posterior probability distribution. We do not find noticeable biases in recovery for any of the other parameters.
@@ V Conclusions
[…]
@@
We also consider the parameter biases for ECOs measured assuming they are BBH. We find that the mass ratio is biassed towards more asymmetric masses. This suggests that, when considering whether any current observations are ECOs, we should focus on those measured to have unequal masses. For instance, an event like GW190814 [67], where we observed a high mass ratio, could be an ECO, but based on fig. 5 must have $C>0.33$ since that the best-fitting BBH template has a fitting factor of 0.970.97[63].
@@
These measurement results highlight a number of caveats to this work. We have included only one physical effect in potential ECOs, namely a simple model of the effect of compactness on the contact point that terminates inspiral. The model is entirely phenomenological and does not make additional checks on the physics of the source. However, the ability of PhenomDECO to smoothly sample over this new parameter and yielding consistent results with real data highlights the framework’s capacity to embed additional source physics within phenomenological models. The framework can be extended to include physics-based consistency checks, such as considerations for energy conservation due to the early end of inspiral, which we leave for future work. Furthermore, not all ECOs will follow this phenomenology. For example, consider one of the boson star simulations in Ref. [42] where, even with a compactness of $C=0.2$, the signal is almost identical to that from a BBH. Aside from ignoring tidal effects and post-merger dynamics, this is also partly because that work uses a different interpretation for compactness than the usual $M/R$ definition used for BHs. This reminds us that our results apply to ECOs where compactness impacts the signal in one specific way, but does not necessarily apply to all ECOs, or all definitions of compactness.
@@
Similarly, a strength of such an approach also is that it allows us to determine the properties of compact objects that could be measured without making any prior assumptions about the nature of the object. By contrast, if one produces a model of signals from, for example, boson-star binaries, and then observes a signal consistent with that model (and nott consistent with BBH), we cannot then conclude that the objects are in fact boson stars: only that they are more like boson stars than black holes. They may in fact be the manifestation of some entirely different physical phenomena that leads to compact objects that, when in binaries, produce GW signals more akin to boson-star binaries than black-hole binaries. A more agnostic approach would allow us to measure general properties of ECOs, and from there determine the range of physical processes that could produce such objects. Once we identify which physical properties have the largest influence on the observed signal morphology, we can prioritize these features in follow-up NR simulations. Therefore, compactness estimates from PhenomDECO can inform NR simulation priorities – and of course, simulation results are necessary to refine PhenomDECO  as an ECO identifier.